# Catastrophic Failure of LLM Unlearning via Quantization

### v2

@@ Appendix H Ablation Study
Appendix I Experiments and discussion on other unlearning methods
@@
Appendix J Further Insights into the Core Contributions
@@
  • However, as demonstrated in the experimental results in the ablation study (Appendix H), a large learning rate can lead to model performance degradation. Thus, we extend the concept of localization-informed unlearning methods to the domain of LLMs by calculating module-level gradients and constructing saliency maps. It is important to note that this is not the only approach to mitigating the negative impact of a large learning rate. Our aim here is to make an initial effort to empirically verify whether localization-informed unlearning could be a promising solution to help mitigate the effects of a large learning rate. Overall, our initial efforts validate the theoretical analysis and inspire future research aimed at increasing weight changes during unlearning to prevent knowledge recovery via quantization. We thank the reviewer for pointing this out and will include a discussion in a future version of our paper.
Appendix K The critical nature and real-world scenarios of the identified problem
@@
Appendix L Comparison with other methods for evaluating robust unlearning in LLMs
@@
  • Our motivation is not to attack the unlearned model and force it to recover the knowledge. Instead, we consider the problem from the perspective that, as quantization becomes increasingly popular in the era of LLMs, it is important to evaluate whether existing unlearning methods remain effective when the unlearned model is quantized.
  • the performance of forcing an unlearned model to generate the supposed forgotten knowledge is poor. Specifically, the unlearned model retains only 10% of the knowledge, and none of the attack methods (including Other Language, Jailbreaks, In-Context Extraction, Fine-Tuning, Downstream Task, Latent Knowledge, Prompting Baselines, Side Effects) achieve the goal of forcing the unlearned model to retain 20% of the knowledge. These results are much less effective compared to our empirical findings. Moreover, the attack methods mentioned in (Lynch et al., 2024), such as Jailbreak and In-context Learning, require significantly more effort to recover knowledge from an unlearned model compared to the simplicity and effectiveness of quantization.
Appendix M Reproducibility
We provide experimental setup and implementation details in Appendix D. Our code is available at: https://anonymous.4open.science/r/FailureUnlearning-20DE.

### v3

@@ Appendix I Ablation Study
Appendix J Experiments and discussion on other unlearning methods
@@
Appendix K Further Insights into the Core Contributions
@@
  • However, as demonstrated in the experimental results in the ablation study (Appendix I), a large learning rate can lead to model performance degradation. Thus, we extend the concept of localization-informed unlearning methods to the domain of LLMs by calculating module-level gradients and constructing saliency maps. It is important to note that this is not the only approach to mitigating the negative impact of a large learning rate. Our aim here is to make an initial effort to empirically verify whether localization-informed unlearning could be a promising solution to help mitigate the effects of a large learning rate. Overall, our initial efforts validate the theoretical analysis and inspire future research aimed at increasing weight changes during unlearning to prevent knowledge recovery via quantization. We thank the reviewer for pointing this out and will include a discussion in a future version of our paper.
Appendix L The critical nature and real-world scenarios of the identified problem
@@
Appendix M Comparison with other methods for evaluating robust unlearning in LLMs
@@
  • Our motivation is different from attacking the unlearned model and forcing it to recover the knowledge. Instead, we consider the problem from the perspective that, as quantization becomes increasingly popular in the era of LLMs, it is important to evaluate whether existing unlearning methods remain effective when the unlearned model is quantized.
  • The performance of forcing an unlearned model to generate the supposed forgotten knowledge is poor (Lynch et al., 2024). Specifically, the unlearned model retains only 10% of the knowledge, and none of the attack methods (including Other Language, Jailbreaks, In-Context Extraction, Fine-Tuning, Downstream Task, Latent Knowledge, Prompting Baselines, Side Effects) achieve the goal of forcing the unlearned model to retain 20% of the knowledge. These results are much less effective compared to our empirical findings. Moreover, the attack methods mentioned in (Lynch et al., 2024), such as Jailbreak and In-context Learning, require significantly more effort to recover knowledge from an unlearned model compared to the simplicity and effectiveness of quantization.
Appendix N Reproducibility
We provide experimental setup and implementation details in Appendix D. Our code is available at: https://github.com/zzwjames/FailureLLMUnlearning.